# LIO-GVM: an Accurate, Tightly-Coupled Lidar-Inertial Odometry with Gaussian Voxel Map

### v1

@@ V-A Evaluation of Odometry Accuracy
The accuracy of odometry is evaluated by comparing the KITTI metric [30], the average translation error (ATE), and the average rotation error (ARE):
$ARE$ $=\sum_{\left(i,j\right)\in\mathcal{F}}rot({}^{G}\bar{\mathbf{T}}_{j}{}^{G}\bar{\mathbf{T}}_{i}^{-1}{}^{G}\mathbf{T}_{j}{}^{G}\mathbf{T}_{i}^{-1}),$ (26)
$ATE$ $=\sum_{\left(i,j\right)\in\mathcal{F}}\lVert{}^{G}\bar{\mathbf{T}}_{j}{}^{G}\bar{\mathbf{T}}_{i}^{-1}{}^{G}\mathbf{T}_{j}{}^{G}\mathbf{T}_{i}^{-1}\rVert_{2},$
where $\mathcal{F}$ is a set of frames $(i,j)$, ${}^{G}\bar{\mathbf{T}}_{j}$ and ${}^{G}\mathbf{T}_{j}$ are the estimated and ground truth poses respectively. $rot(\cdot)$ is the rotation angle. We compare LIO-GVM with three state-of-the-art LiDAR and LiDAR-IMU odometry algorithms: DLO [31], LIO-SAM [1], and FAST-LIO2 [10]. DLO is a GICP-based LO while LIO-SAM and FAST-LIO2 are LIO systems. It should be noted that no loop closures have been performed for all the methods.
@@
Two variations of the proposed system LIO-GVM(w/o s) and LIO-GVM(plane), are tested for ablation study. LIO-GVM(w/o s) is implemented without the outlier rejecting correspondence matching scheme proposed in Sec. IV-C1; the similarity is still utilized in the residual formulation. LIO-GVM(plane) replaces the proposed new residual metric with the point-to-plane distance [10]. As shown in Tab. II, both LIO-GVM(w/o s) and LIO-GVM(plane) turn worse compared to the original system for all the sequences, representing that the proposed correspondence matching scheme and residual metric both contribute to the accuracy of the system. The accuracy of LIO-GVM(w/o s) tends to be much worse than LIO-GVM(plane), which is because the residual metric inherently reduces the weights for mismatches. The deterioration of the odometry is particularly evident for the ntuo_* sequences, in which the robot runs in an aggressive motion and is equipped with the low-cost IMU.
As illustrated in Tab. II, LIO-GVM reaches the best results for the most sequences. Note that LIO-SAM is not applicable for ntuo_* and nc_ sequences as it requires the equipment of a 9-axis IMU. For the ntu_* sequences, LIO-GVM achieves comparable accuracy to FAST-LIO2, outperforming the other algorithms. It is observed that LIO-GVM shows obvious superiority on ntuo_* sequences, which illustrates that when equipped with low-cost IMU, LIO-GVM can achieve better performance compared with other LIO systems. Similarly, the results of LIO-GVM on nc_01 and nc_02 demonstrate its ability to deal with accumulated drifts.
@@
Overall, LIO-GVM has demonstrated the best performance regarding odometry accuracy, showing the effectiveness of the error-rejecting correspondence matching scheme and the distribution-based residual formulation in achieving precise location estimation across various sensor setups, especially in long-term applications or with cheaper IMUs.
ntu_02
ntu_04
ntu_10
ntu_13
ntuo_02
ntuo_04
ntuo_10
ntuo_13
nc_01
nc_02
nc_05
nc_06
nc_07
DLO
6.46/0.43
3.65/0.11
7.65/0.51
3.98/0.20
6.46/0.43
3.75/0.11
7.65/0.51
3.98/0.20
4.20/0.51
4.27/0.37
3.44/0.44
1.31/0.62
1.73/0.18
LIO-SAM
7.80/0.52
3.71/0.23
9.32/0.63
3.05/0.16
-${}^{a}$
FAST-LIO2
2.09/0.24
2.11/0.13
2.34/0.15
2.89/0.11
5.64/0.80
3.51/0.23
7.32/0.41
6.79/0.50
3.23/0.37
4.15/0.43
3.13/0.26
0.94/0.37
1.16/0.15
LIO-GVM(w/o s)
2.37/0.40
2.35/0.11
3.21/0.14
2.79/0.14
5.51/0.68
4.73/0.39
6.59/0.39
4.58/0.26
3.36/0.31
4.76/0.44
3.46/0.47
1.28/0.43
1.38/0.20
LIO-GVM(plane)
8.24/0.68
6.91/0.41
8.29/0.74
5.77/0.43
9.34/0.90
7.95/0.54
10.2/0.81
8.89/0.73
6.29/0.49
6.30/0.67
4.54/0.72
2.48/0.59
3.45/0.30
LIO-GVM
1.74/0.46
2.26/0.05
3.16/0.11
2.74/0.09
3.23/0.40
2.69/0.14
4.25/0.27
3.60/0.18
2.47/0.26
3.20/0.35
3.27/0.33
0.89/0.33
1.10/0.17
  • All the errors are represented as ATE[%]/ARE[deg/10m] (the smaller the better).
  • Bold value stands for the best value; underline value stands for the second best value; ’-’ denotes that the algorithm is not applicable.
  • ${}^{a}$ The built-in IMU of OS1-128 and OS1-64 cannot provide attitude data, which is compulsory for LIO-SAM.
V-B Evaluation of Storage Efficiency
TABLE III: The constructed Map Storage analysis
ntu_02
ntu_04
ntu_10
ntu_13
nc_01
nc_02
nc_05
nc_06
nc_07
FAST-LIO2
6.07
21.83
19.13
22.19
7.93
9.24
1.15
0.65
5.62
LIO-GVM
3.55 12.61 11.39 8.29 4.25 4.60 0.13 0.48 3.01
  • The map storage is analyzed by evaluating the storage consumption [MB] for saving the eventually constructed map.
  • Bold value stands for the best value.
In this section, we evaluate the storage consumption of the global map generated by LIO-GVM. We employ the same parameters as those used in the experiments for evaluating the odometry accuracy. Specifically, in LIO-GVM, we have defined a custom point type using the Point Cloud Library (PCL). Each point encapsulates all the elements of $(\boldsymbol{\mu}_{i},\mathbf{C}_{i},M)$, comprising a total of four integer values and nine floating-point values. We save all the voxels from the global map into a .pcd (Point Cloud Data) file format. For the other systems under comparison, we configure their local map size to be 500m, ensuring coverage of all areas within the two datasets, and save all the points of the map to a .pcd file using the same function.
As illustrated in Tab. III, the map storage consumption of LIO-GVM outperforms other systems for all the sequences. This is because other systems need to ensure the downsampling filter of the map is not excessively large (usually $<$ 0.5m) to preserve the geometric information of the points. In contrast, our approach, which involves maintaining merely four integer values and nine floating-point values for each voxel yet still offering a comprehensive representation of the intra-voxel point distribution, permits us to allocate a considerably larger voxel size (commonly $>$ 1m) compared to other systems. Considering that all the LiDAR scans from the 11 sequences are downsampled as a pre-processing step in all the evaluated systems, the extent of improvement in map storage of LIO-GVM compared to other systems may not exhibit significant variations for different LiDAR resolutions.
V-C Evaluation of Time Consumption
In this section, we will evaluate the temporal efficiency of LIO-GVM, utilizing the same parameters as those deployed in our previous odometry evaluations for consistency. We first analyze the influence of the data structure on the mapping procedure, which involves correspondence matching and incremental updates. Given the superior performance of the ikd-Tree provided in [10] compared to other dynamic tree-based structures, our comparison will be limited to this structure. While the Gaussian fitting process does not directly participate in the mapping procedure, its time consumption is still accounted for in LIO-GVM. This is to ensure a fair comparison, as LIO-GVM requires that the points from the input scan are modeled as a Gaussian distribution, thereby indirectly involving the Gaussian fitting. We denote the data structure we use in LIO-GVM for the mapping procedure as GVM. The results are illustrated in Tab. IV. For Gaussian Fitting, ikd-Tree exhibits no time consumption, as this process is not required for this data structure. GVM significantly outperforms ikd-Tree in correspondence matching because the time complexity of $k$-nearest neighbors searching for ikd-Tree is $\mathcal{O}(log(n))$, where $n$ is the tree size; but for GVM, managed in a hash table, the corresponding time complexity is $\mathcal{O}(1)$, which is substantially faster than ikd-Tree. ikd-Tree slightly performs better in the incremental update. Nonetheless, GVM demonstrates superior performance in total map time.
In Tab. V, we illustrate the average total time consumed per scan for all the sequences. The total time we evaluated includes all the procedures: data process, pose estimate, and map update. LIO-GVM and FAST-LIO2 perform map update every time a pose is estimated. But for LIO-SAM and DLO, they run map update in another thread with the pose estimation thread; thus we only evaluate their time consumed for state estimation. Even though the mapping procedure of LIO-GVM is faster than FAST-LIO2, FAST-LIO2 achieves the shortest total time for almost all the sequences, followed by DLO, the second best. One possible reason for this could be that LIO-GVM requires more geometric information from the input scan, leading to a higher-dimensional measurement matrix. However, as illustrated in Tab. VI, the fast performance of FAST-LIO2 and DLO is at the cost of higher runtime RAM usage.
TABLE IV: evaluation of the temporal efficiency in sequences on mapping procedures
@@ V-C Evaluation of Time Consumption
ikd-Tree
GVM
ikd-Tree
GVM
ikd-Tree
GVM
ikd-Tree
GVM
ntu_02
0
11.61
25.05
1.29 2.45
5.13
27.50
18.03
ntu_04
0
12.91
24.26
1.41 3.96
5.34
28.22
19.66
ntu_10
0
11.86
24.27
1.34 3.97
5.87
28.24
19.08
ntu_13
0
12.71
24.89
1.67 3.62
5.18
28.51
19.56
nc_01
0
8.10
14.36
0.69 0.78
2.97
15.14
11.76
nc_02
0
8.25
15.76
0.73 0.92
2.84
16.68
11.82
nc_05
0
7.73
12.62
0.67 0.57
2.84
13.19
11.24
nc_06
0
7.19
13.66
0.70 0.81
3.23
14.47
11.12
nc_07
0
8.99
15.67
0.62 1.11
3.18
16.78
12.79
@@
TABLE V: evaluation of the temporal efficiency in sequences on mapping procedures
ntu_02
ntu_04
ntu_10
ntu_13
nc_01
nc_02
nc_05
nc_06
nc_07
DLO(odom)
39.94
43.63
39.14
40.16
18.35
23,48
20.30
23.64
28.83
LIO-SAM(odom)
70.41
71.11
67.85
70.72
FAST-LIO2 37.64 39.23 38.23 38.51
19.33
21.04 17.21 19.59 20.94
LIO-GVM
47.65
46.91
49.23
47.19
26.77
28.33
28.51
27.82
30.13
  • the average time consumed per scan [ms] is evaluated.
  • Bold value stands for the best value;
TABLE VI: comparison of rumtime RAM in sequences
ntu_02
ntu_04
ntu_10
ntu_13
nc_01
nc_02
nc_05
nc_06
nc_07
DLO
1112
1564
1727
1389
1102
1438
513
337
809
FAST-LIO2
404
601
738
444
479
610
236
221
361
LIO-GVM
361 529 690 373 450 584 199 191 334
  • The eventually runtime RAM usage [MB] is compared.
  • Bold value stands for the best value;
@@ VI Conclusion
In conclusion, this letter proposes an accurate filter-based LIO system denoted as LIO-GVM, which models the points as Gaussian distributions for robust correspondence matching. A key innovation that distinguishes our work from others is the proposed new residual metric for the filter-based LIO system, which marks an improvement from merely quantifying distance to incorporating variance disparity. Benefiting from the new metric, the map of LIO-GVM can be maintained in a simple yet efficient manner. We demonstrate the reliability and efficiency of LIO-GVM through extensive experiments on different datasets collected in large-scale, perceptually challenging environments. For future works, we would like to develop an adaptive voxel representation scheme for the global map, which can greatly reduce the map size as well as increase the odometry accuracy. Besides, the occupancy information is not utilized for the current work; we plan to design an efficient ray tracing method to verify the occupancy status, enabling applications such as collision detection and path planning.
@@
We provide the detail formulation of the involved matrices in the supplemental material: supplemental_material.pdf.
@@
$\|\mathbf{z}_{k}^{n}\|^{2}_{\mathbf{V}_{k}^{-1}}$ $=\sum_{j=0}^{m}(s_{j,i}^{n})^{2}{}^{G}\tilde{\boldsymbol{\mu}}^{T}(\hat{\mathbf{C}}_{j}^{n}+\mathbf{C}_{i})^{-1}{}^{G}\tilde{\boldsymbol{\mu}}$ (27)
$=\sum_{j=0}^{m}(s_{j,i}^{n})^{2}{}^{G}\tilde{\boldsymbol{\mu}}^{T}\left(\mathbf{U}_{j}^{n}\right)^{T}\left(\boldsymbol{\Lambda}_{j}^{n}\right)^{-1}\mathbf{U}_{j}^{n}{}^{G}\tilde{\boldsymbol{\mu}}$
$=\sum_{j=0}^{m}(s_{j,i}^{n})^{2}{}^{G}\tilde{\boldsymbol{\mu}}^{T}\left(\mathbf{U}_{j}^{n}\right)^{T}\left(\mathbf{Q}_{j}^{n}\right)^{T}V_{k}^{-1}\mathbf{Q}_{j}^{n}\mathbf{U}_{j}^{n}{}^{G}\tilde{\boldsymbol{\mu}}$
$=\sum_{j=0}^{m}s_{j,i}^{n}{}^{G}\tilde{\boldsymbol{\mu}}^{T}\left(\mathbf{U}_{j}^{n}\right)^{T}\left(\mathbf{Q}_{j}^{n}\right)^{T}V_{k}^{-1}s_{j,i}^{n}\mathbf{Q}_{j}^{n}\mathbf{U}_{j}^{n}{}^{G}\tilde{\boldsymbol{\mu}}$
$=\sum_{j=0}^{m}(\mathbf{z}_{k,j}^{n})^{T}V_{k}^{-1}\mathbf{z}_{k,j}^{n},$
@@
-C Derivation of Jacobian Matrix $H_{j}$
The Jacobian matrix would be:
$H_{j}^{n}=\frac{\partial{\mathbf{z}_{j}(\hat{\mathbf{x}}_{k}\boxplus\delta\mathbf{x}_{k}^{n},{}^{L_{k}}\mathbf{p}_{j})}}{\partial{\delta\mathbf{x}_{k}^{n}}}$ (29)
the observation function is:
$\mathbf{z}_{j}(\hat{\mathbf{x}}_{k}\boxplus\delta\mathbf{x}_{k}^{n},{}^{L}\mathbf{p}_{j})={s_{j,i}^{n}}{\mathbf{D}_{j}}^{n}\left({}^{G}\hat{\mathbf{p}}_{j}-{}^{G}\boldsymbol{\mu}_{i}\right)$ (30)
The calculation of the derivative of $s_{j,i}^{n}\mathbf{D}_{j}^{n}$ w.r.t. $\delta\mathbf{x}_{k}^{n}$ is quite comprehensive and impractical. We assume $s_{j,i}^{n}\mathbf{D}_{j}^{n}$ as constant for each iteration and update its value after the iteration in the implementation. According to the experiment results, this assumption can provide enough accuracy while greatly reducing the computation load. Let’s pre-process the observation function:
$\mathbf{z}_{j}(\hat{\mathbf{x}}_{k}\boxplus\delta\mathbf{x}_{k}^{n},{}^{L}\mathbf{p}_{j})={s_{j}}\mathbf{D}_{j}^{n}\biggl{(}{}^{G}\hat{\mathbf{p}}_{I_{k}}+{}^{G}\delta\mathbf{p}_{I_{k}}-{}^{G}\boldsymbol{\mu}_{j}$ (31)
$+{}^{G}\hat{\mathbf{R}}_{I_{k}}\mathrm{Exp}\left({}^{G}\delta\mathbf{r}_{I_{k}}\right)\left({}^{I}{\mathbf{R}}_{L_{k}}{}^{L}\mathbf{p}_{j}+{}^{I}{\mathbf{p}_{L_{k}}}\right)\biggr{)}.$
Due to the minute transformation between two LiDAR frame, we have:
$\mathrm{Exp}\left({}^{G}\delta\mathbf{r}_{I_{k}}\right)\approx\mathbf{I}+\left[{}^{G}\delta\mathbf{r}_{I_{k}}\right]_{\times}.$ (32)
Combine (31) and (32) we have:
$\mathbf{z}_{j}(\hat{\mathbf{x}}_{k}\boxplus\delta\mathbf{x}_{k}^{n},{}^{L}\mathbf{p}_{j})={s_{j}}\mathbf{D}_{j}^{n}\biggl{(}{}^{G}\hat{\mathbf{p}}_{I_{k}}+{}^{G}\delta\mathbf{p}_{I_{k}}-{}^{G}\boldsymbol{\mu}_{j}$ (33)
$+{}^{G}\hat{\mathbf{R}}_{I_{k}}{}^{I}\mathbf{q}_{j}-{}^{G}\hat{\mathbf{R}}_{I_{k}}\left[{}^{I}\mathbf{q}_{j}\right]_{\times}{}^{G}\delta\mathbf{r}_{I_{k}}\biggr{)},$
where ${}^{I}\mathbf{q}_{j}={}^{I}{\mathbf{R}}_{L_{k}}{}^{L}\mathbf{p}_{j}+{}^{I}{\mathbf{p}_{L_{k}}}$. It is easy to derive the Jacobian matrix $\mathbf{H}_{j}^{n}$ as following:
$\mathbf{H}_{j}^{n}=\begin{bmatrix}-{s_{j}}\mathbf{D}_{j}^{n}\left[{}^{I}\mathbf{q}_{j}\right]_{\times}&{s_{j}}\mathbf{D}_{j}^{n}&\mathbf{0}_{3\times 15}\end{bmatrix}$ (34)

### v3

@@ V-A Evaluation of Odometry Accuracy
We evaluate the accuracy by comparing the KITTI metric [28]: the average translation error (ATE), and the average rotation error (ARE). We compare LIO-GVM with three state-of-the-art LiDAR (-inertial) odometry algorithms: DLO [29], VoxelMap [12], LIO-SAM [1], and FAST-LIO2 [4]. DLO is a GICP-based LO, VoxelMap is a probabilistic voxel-based LO while LIO-SAM and FAST-LIO2 are LIO systems. Parameters for all methods are kept constant across sequences for consistency. For LIO-GVM, the voxel size, $s_{t}$, and $\alpha$ are empirically set as 1.0 $m$, 0.70 and $10^{-6}$, respectively. Additionally, the neighbor searching scheme is set as 7-neighbors searching.
Figure 5: The position error of different methods on sequence ntu_13. We find that the position error is most significant along the z axis, due to the severe elevation changes.
Figure 6: The trajectories reference to the ground on sequence ntu_04.
@@
Two variations of the proposed system denoted as LIO-GVM(w/o s) and LIO-GVM(plane) are tested for ablation study. LIO-GVM(w/o s) is implemented without the outlier rejecting scheme proposed in Sec. IV-C1. LIO-GVM(plane) replaces the proposed new residual metric with the point-to-plane distance [4]. As shown in Tab. II, both LIO-GVM(w/o s) and LIO-GVM(plane) turn worse compared to the original system for all the sequences, indicating that the proposed correspondence matching scheme and residual metric both contribute to the accuracy of the system. The accuracy of LIO-GVM(plane) tends to be much worse than LIO-GVM(w/o s), which is because the residual metric inherently reduces the weights for mismatches. The drastic accuracy degradation of LIO-GVM(plane) is because, without the proposed residual metric, all the adopted correspondences contribute equally to the estimation, even those with low similarity.
As illustrated in Tab. II, LIO-GVM achieves the best results for most sequences due to its ability to reject mismatched distributions (the proposed correspondence matching) and evaluate the variance divergence in the distribution covariance (the proposed residual metric). For ntu_* sequences, FAST-LIO2 achieves comparable results with LIO-GVM, because the accurate IMU integration can approximately project source points to the correct feature in the target map. VoxelMap offers accurate pose estimation in some sequences but frequently diverges, especially for the nc_* sequences equipped with a 64-line LiDAR. This divergence is attributed to its map update scheme. When the LiDAR remains stationary for several scans, distant voxels converge prematurely and stay undersampled. This issue is particularly obvious for ground voxels, leading to the majority of divergence occurrences along the z-axis. Conversely, LIO-GVM employs a continuously updated mapping scheme, ensuring that the performance remains stable for motion and LiDAR sampling rate variations.
The performance of all LIO systems is negatively impacted in ntuo_* compared to ntu_* sequences due to the equipment of a low-cost IMU. Nevertheless, LIO-GVM is able to mitigate the impact of IMU noises through its proposed correspondence matching scheme, which can reject outliers. Additionally, LIO-GVM incorporates a new residual metric that takes similarity into account during the estimation phase, allowing more effective correspondences to have a greater influence on the estimation. Consequently, LIO-GVM exhibits the best accuracy compared to other systems. However, all the methods have a common limitation in accommodating the pronounced changes in elevation in the ntu_* sequences, as seen in Fig. 5 and Fig. 6.
@@
Overall, LIO-GVM has demonstrated the best performance regarding odometry accuracy, showing the effectiveness of the error-rejecting correspondence matching scheme and the distribution-based residual metric in achieving accurate pose estimation across various sensor setups, especially in long-term applications or with cheaper IMUs.
TABLE III: evaluation of the temporal efficiency in sequences on mapping procedures
@@
ikd-Tree GVM ikd-Tree GVM ikd-Tree GVM ikd-Tree GVM
ntu_02 0 11.61 25.05 1.29 2.45 5.13 27.50 18.03
ntu_04 0 12.91 24.26 1.41 3.96 5.34 28.22 19.66
ntu_10 0 11.86 24.27 1.34 3.97 5.87 28.24 19.08
ntu_13 0 12.71 24.89 1.67 3.62 5.18 28.51 19.56
nc_01 0 8.10 14.36 0.69 0.78 2.97 15.14 11.76
nc_02 0 8.25 15.76 0.73 0.92 2.84 16.68 11.82
nc_05 0 7.73 12.62 0.67 0.57 2.84 13.19 11.24
nc_06 0 7.19 13.66 0.70 0.81 3.23 14.47 11.12
nc_07 0 8.99 15.67 0.62 1.11 3.18 16.78 12.79
@@
(a) MCD_VIRAL dataset.
(b) Newer College dataset.
Figure 7: The average time consumed per scan (ms) for all methods in different datasets.
V-B Evaluation of Storage Efficiency
In this section, we evaluate the storage consumption of the global map generated by LIO-GVM. Specifically, in LIO-GVM, we have defined a custom point type using the Point Cloud Library (PCL). Each point encapsulates all the elements of $(\boldsymbol{\mu}_{i},\mathbf{C}_{i},M)$, comprising a total of four integer values and nine floating-point values. All the voxels are stored in a compressed binary .pcd file. For other systems, we configure their local map size to be 500m, ensuring coverage of all areas within the two datasets, and save all the points in the map to a compressed binary .pcd file.
Figure 8: The eventual runtime RAM usage for all the methods.
As illustrated in Tab. IV, the map storage consumption of LIO-GVM outperforms other systems for all the sequences. This is because other systems need to ensure that the downsampling filter of the map is not excessively large (usually $<$ 0.5m) to preserve the geometric information of the points. In contrast, our approach allows us to allocate a considerably larger voxel size (commonly $>$ 1m) compared to other systems while still offering a comprehensive representation of the intra-voxel point distribution.
TABLE IV: The constructed Map Storage analysis
ntu_02 ntu_04 ntu_10 ntu_13 nc_01 nc_02 nc_05 nc_06 nc_07
FAST-LIO2 6.07 21.83 19.13 22.19 7.93 9.24 1.15 0.65 5.62
LIO-GVM 3.55 12.61 11.39 8.29 4.25 4.60 0.13 0.48 3.01
  • The map storage is analyzed by evaluating the storage consumption [MB]. All the maps are saved as binary compressed PCD files.
  • Bold value stands for the best value.
V-C Evaluation of Time Consumption
In this section, we will evaluate the temporal efficiency of LIO-GVM, utilizing the same parameters as those deployed in our previous odometry evaluations for consistency. We first analyze the influence of the data structure on the mapping procedure, which involves correspondence matching and incremental updates. Given the superior performance of the ikd-Tree provided in [4] compared to other dynamic tree-based structures, our comparison will be limited to this structure. While the Gaussian fitting process does not directly participate in the mapping procedure, its time consumption is still accounted for in LIO-GVM for a fair comparison. We denote the data structure in LIO-GVM for the mapping procedure as GVM. The results are illustrated in Tab. III. GVM significantly outperforms ikd-Tree in correspondence matching because the time complexity of $k$-nearest neighbors searching for ikd-Tree is $\mathcal{O}(log(n))$, where $n$ is the tree size; but for GVM, managed in a hash table, the corresponding time complexity is $\mathcal{O}(1)$. ikd-Tree slightly performs better in the incremental update. Nonetheless, GVM demonstrates superior performance in total map time.
In Fig. 7, we illustrate the average total time consumed per scan for all the sequences. The total time we evaluated includes all the procedures: data processing, pose estimation, and map update. Since LIO-SAM and DLO run map update in different thread with the pose estimation thread, we only evaluate their time consumed for state estimation. Notably, VoxelMap consumes an average of 141.3 ms per scan due to its demand for dense point cloud data. Even though the mapping procedure of LIO-GVM is faster than FAST-LIO2, FAST-LIO2 achieves the shortest total time for almost all the sequences. This is because LIO-GVM requires more geometric information from the input scan. For example, in ntu_* sequences, FAST-LIO2 employs fewer feature points than LIO-GVM (4638 versus 5741). Consequently, LIO-GVM spends more time on pose estimation. However, as illustrated in Tab. 8, the fast performance of FAST-LIO2 and DLO comes with the trade-off of increased runtime RAM usage, while LIO-GVM can achieve comparable temporal efficiency with lower RAM usage.
@@ VI Conclusion
In conclusion, this letter proposes an accurate filter-based LIO system denoted as LIO-GVM, which models the points as Gaussian distributions for robust correspondence matching. The key innovation of the system is the proposed new residual metric. Unlike previous methods that solely measure distance, our approach incorporates variance disparity, thereby enhancing the system’s performance. Benefiting from the new metric, the map of LIO-GVM can be maintained in a simple yet efficient manner. We demonstrate the reliability and efficiency of LIO-GVM through extensive experiments on different datasets. The limitation of LIO-GVM is that it requires equipment with higher resolution LiDAR to ensure informative Gaussian fitting of the input scan. Additionally, the voxel size is not adaptive to the environment. To address these limitations, future work will focus on developing an adaptive voxel representation scheme for the global map.
@@
Please find details in the supplemental material: supplemental_material.pdf.
@@
$\|\mathbf{z}_{k}^{n}\|^{2}_{\mathbf{V}_{k}^{-1}}=\sum_{j=0}^{m}(s_{j,i}^{n})^{2}{}^{G}\tilde{\boldsymbol{\mu}}^{T}(\hat{\mathbf{C}}_{j}^{n}+\mathbf{C}_{i}+\alpha\mathbf{I})^{-1}{}^{G}\tilde{\boldsymbol{\mu}}.$ (19)
To avoid the singularity issue, we first perform eigendecomposition on $(\hat{\mathbf{C}}_{j}^{n}+\mathbf{C}_{i}+\alpha\mathbf{I})^{-1}$ to have $(\hat{\mathbf{C}}_{j}^{n}+\mathbf{C}_{i}+\alpha\mathbf{I})^{-1}=\left(\mathbf{U}_{j}^{n}\right)^{T}\left(\boldsymbol{\Lambda}_{j}^{n}\right)^{-1}\mathbf{U}_{j}^{n}$. Subsequently, the eigenvalues are normalized and clamped at $10^{-4}$: $\lambda_{t,j}^{n}=max(\lambda_{t,j}^{n}/trace({\Lambda}_{j}^{n}),10^{-4}),t=1,2,3$. In order to obtain the same formula as the left-hand side, we perform matrix decomposition to have $\left({\Lambda}_{j}^{n}\right)^{-1}=\left(\mathbf{Q}_{j}^{n}\right)^{T}\mathbf{V}_{k}^{-1}\mathbf{Q}_{j}^{n}$, which is quite easy since $\left({\Lambda}_{j}^{n}\right)^{-1}$ is a diagonal matrix. Thus,
[…]
$=\sum_{j=0}^{m}s_{j,i}^{n}{}^{G}\tilde{\boldsymbol{\mu}}^{T}\left(\mathbf{U}_{j}^{n}\right)^{T}\left(\mathbf{Q}_{j}^{n}\right)^{T}\mathbf{V}_{k}^{-1}s_{j,i}^{n}\mathbf{Q}_{j}^{n}\mathbf{U}_{j}^{n}{}^{G}\tilde{\boldsymbol{\mu}}$
$=\sum_{j=0}^{m}(\mathbf{z}_{k,j}^{n})^{T}\mathbf{V}_{k}^{-1}\mathbf{z}_{k,j}^{n},$